# Generalized Moving Least-Squares for Solving Vector-valued PDEs on Unknown Manifolds

### v1

@@ 3 GMLS using intrinsic formulation
In this section, we approximate differential operators for vector fields using intrinsic formulation. In Section 3.1, we review GMLS for the local Monge patch parametrization of each stencil $S_{{\mathbf{x}_{0}}}$ [27, 20]. In Section 3.2, we derive the numerical scheme to approximate the Bochner Laplacian acting on a vector field. Then we approximate the covariant derivative in Section 3.3. For readers not familiar with intrinsic differential geometry, please refer to Appendix A.1 for a short review.
@@ 3.1 Monge parametrization
Then the “height” $q$ is approximated as a polynomial function of the local coordinates $(\theta_{1},\theta_{2})$ using the GMLS approach following the similar procedure introduced in Section 2.2. Here one key difference is the degrees of all “intrinsic” polynomials in the hypothesis space ranging from $2$ to $l$, that is, the hypothesis space is $\mathbb{P}_{\mathbf{x}_{0}}^{l}\backslash\mathbb{P}_{\mathbf{x}_{0}}^{1}=\mathrm{span}(\{\theta_{1}^{\alpha_{1}}\theta_{2}^{\alpha_{2}}\}_{2\leq\alpha_{1}+\alpha_{2}\leq l})$. Thus, the approximation of $q$ is the optimal solution to the least-squares problem (2):
@@
The local Riemannian metric $g=(g_{ij})_{i,j=1}^{2}$ for ${\mathbf{x}}\in S_{{\mathbf{x}_{0}}}$ is approximated as
$g=\left[\begin{matrix}1+\left(\partial_{\theta_{1}}\hat{q}\right)^{2}&\partial_{\theta_{1}}\hat{q}\partial_{\theta_{2}}\hat{q}\\ \partial_{\theta_{1}}\hat{q}\partial_{\theta_{2}}\hat{q}&1+\left(\partial_{\theta_{2}}\hat{q}\right)^{2}\end{matrix}\right].$ (11)
As usual, we use $g^{ij}$ to denote the inverse metric tensor. Other geometric quantities can be similarly computed in the stencil $S_{{\mathbf{x}_{0}}}$.
@@ 3.2 Approximation of the Bochner Laplacian
For any vector field given on the point cloud $\mathbf{X}_{M}$, we write it as $u=u^{j}\bm{t}_{j}$ where the components $\{u^{1},u^{2}\}$ are defined on the nodes $\{\mathbf{x}_{i}\}_{i=1}^{N}$ and $\left\{\bm{t}_{1}(\mathbf{x}_{i}),\bm{t}_{2}(\mathbf{x}_{i})\right\}_{i=1}^{N}$ are bases of tangent vector space at each node. Note that there is no need to do alignments of these bases (as shown in Fig. 1(a)). To obtain a finite-difference type approximation of the Bochner Laplacian at a base point $\mathbf{x}_{0}$, we intend to find the weights $\{\bm{w}^{k}\in\mathbb{R}^{2\times 2}\}_{k=1}^{K}$ such that,
$\Delta_{B}u({\mathbf{x}_{0}})\approx\sum_{k=1}^{K}\bm{w}^{k}u({\mathbf{x}_{0,k}})=(\bm{w}^{1},\ldots,\bm{w}^{K})\mathbf{u}_{{\mathbf{x}_{0}}},$ (12)
where $\{{\mathbf{x}_{0,k}}\}_{k=1}^{K}$ is the $K$-nearest neighbors of $\mathbf{x}_{0}$, $u({\mathbf{x}_{0,k}})=(u^{1}({\mathbf{x}_{0,k}}),u^{2}({\mathbf{x}_{0,k}}))^{\top}$ and $\ \mathbf{u}_{{\mathbf{x}_{0}}}=(u^{1}({\mathbf{x}_{0,1}}),u^{2}({\mathbf{x}_{0,1}}),\ldots,u^{1}({\mathbf{x}_{0,K}}),u^{2}({\mathbf{x}_{0,K}}))^{\top}\in\mathbb{R}^{2K\times 1}.$
In Section 3.2.1, we derive the intrinsic formulation for the Bochner Laplacian in the local Monge coordinate system (10). In Section 3.2.2, we apply the GMLS to discretize the operator in the local Monge coordinate. Notice that the representation of vector fields in the local coordinate chart is different from that in the global $\{\bm{t}_{1},\bm{t}_{2}\}$. In Section 3.2.3, we discuss how to obtain the weights in (12) and assemble these weights into a sparse $2N\times 2N$ matrix where $2$ corresponds to the intrinsic dimension of the surface.
@@ 3.2.1 Approximation of geometric quantities in the local Monge coordinate
The computations for geometric quantities and function derivatives involved in the Bochner Laplacian $\Delta_{B}u$ could be complicated. Fortunately, here we only focus on the values of $\Delta_{B}u$ at the base point ${\mathbf{x}_{0}}$ and their calculations can be much simplified by noticing that the Monge coordinate system as introduced in Section 3.1 is a normal coordinate system. In particular, we use the following properties of a normal coordinate system:
@@
Let $\{U_{0};\xi^{1},\ldots,\xi^{d}\}$ be a normal coordinate system at $\mathbf{x}_{0}$ where $U_{0}\subset M$ is an open neighborhood of $\mathbf{x}_{0}$ and $M$ is a $d$-dimensional manifold with Riemannian metric $g$. Denote the Christoffel symbol by $\Gamma_{ij}^{k}$. Then
@@
The Bochner laplacian acting on a vector field $u=\tilde{u}^{i}\partial_{i}$ is defined as
$\Delta_{B}u=g^{ij}\tilde{u}^{k}_{;ij}\partial_{k},$
@@
Applying Proposition 3.1 and noticing that the tangent plane bases in (10) are $\partial_{k}|_{\mathbf{x}_{0}}=\bm{t}_{k}({\mathbf{x}_{0}})$, one can simplify the Bochner Laplacian $\Delta_{B}u(\mathbf{x}_{0})$ as,
$\Delta_{B}u(\mathbf{x}_{0})=g^{ij}\tilde{u}_{;ij}^{k}\partial_{k}|_{\mathbf{x}_{0}}=\left[\left(\tilde{u}_{;11}^{1}+\tilde{u}_{;22}^{1}\right)\partial_{1}+\left(\tilde{u}_{;11}^{2}+\tilde{u}_{;22}^{2}\right)\partial_{2}\right]\big{|}_{\mathbf{x}_{0}},$
@@
where $a_{\alpha_{1},\alpha_{2}}$ are the coefficients of $\hat{q}$ as in (9). Similarly, one can obtain all the derivatives of Christoffel symbols at $\mathbf{x}_{0}$ for 2D manifolds:
@@
where the restriction notations $|_{\mathbf{x}_{0}}$ are all omitted. Then one can calculate $\Delta_{B}u(\mathbf{x}_{0})$ in (3.2.1) as
$\Delta_{B}u(\mathbf{x}_{0})$ $=$ $\left(\partial_{11}\tilde{u}^{1}+\partial_{22}\tilde{u}^{1}+\tilde{u}^{1}\left[4(a_{2,0})^{2}+(a_{1,1})^{2}\right]+\tilde{u}^{2}\left[2a_{2,0}a_{1,1}+2a_{0,2}a_{1,1}\right]\right)\bm{t}_{1}({\mathbf{x}_{0}})$ (15)
@@ 3.2.2 Approximation of function and its derivatives in the local Monge coordinate
Let $u=\tilde{u}^{i}\partial_{i}$ be a vector field defined locally on the stencil $S_{\mathbf{x}_{0}}$, where $\left\{\partial_{1},\partial_{2}\right\}$ is the local basis defined in (10). Our goal here is to approximate the Bochner Laplacian $\Delta_{B}$ acting on the vector field $u$ at ${\mathbf{x}_{0}}$ by a linear combination of the vector values on $S_{\mathbf{x}_{0}}$, i.e., $\{\tilde{u}({\mathbf{x}_{0,k}})=(\tilde{u}^{1}({\mathbf{x}_{0,k}}),\tilde{u}^{2}({\mathbf{x}_{0,k}}))^{\top}\in\mathbb{R}^{2\times 1}\}_{k=1}^{K}$. Thus, we are seeking the weights $\{\bm{\tilde{w}}^{k}\in\mathbb{R}^{2\times 2}\}_{k=1}^{K}$ such that,
$\Delta_{B}u({\mathbf{x}_{0}})\approx\sum_{k=1}^{K}\bm{\tilde{w}}^{k}\tilde{u}({\mathbf{x}_{0,k}})=(\bm{\tilde{w}}^{1},\ldots,\bm{\tilde{w}}^{K})\mathbf{\tilde{u}}_{{\mathbf{x}_{0}}}\in\mathbb{R}^{2\times 1},$ (16)
@@
The GMLS approximation of the vector field $u=\tilde{u}^{i}\partial_{i}$ gives the GMLS solution to each component, i.e.,
@@
Substituting all above into $\Delta_{B}u(\mathbf{x}_{0})$ in (15), one arrives at
$\Delta_{B}u(\mathbf{x}_{0})=(\Delta_{B}u(\mathbf{x}_{0}))_{1}\bm{t}_{1}({\mathbf{x}_{0}})+(\Delta_{B}u(\mathbf{x}_{0}))_{2}\bm{t}_{2}({\mathbf{x}_{0}})$
@@
where the first component of $\Delta_{B}u(\mathbf{x}_{0})$ is $(\Delta_{B}u(\mathbf{x}_{0}))_{1}:=\bm{A}_{11}\bm{\beta}+\bm{A}_{12}\bm{\gamma}$ and the second component is $(\Delta_{B}u(\mathbf{x}_{0}))_{2}:=\bm{A}_{21}\bm{\beta}+\bm{A}_{22}\bm{\gamma}$. Here $\bm{\beta}=\left(\beta_{\alpha_{1}(j),\alpha_{2}(j)}\right)_{j=1}^{m}\in\mathbb{R}^{m\times 1}$ and $\bm{\gamma}=\left(\gamma_{\alpha_{1}(j),\alpha_{2}(j)}\right)_{j=1}^{m}\in\mathbb{R}^{m\times 1}$ are both $m$ by $1$ column vectors and $\{\bm{A}_{ij}\in\mathbb{R}^{1\times m}\}_{i,j=1}^{2}$ are all $1$ by $m$ row vectors. Specifically, $\bm{A}_{11}$ has three nonzero entries $24(a_{2,0})^{2}+(a_{1,1})^{2},2,$ 2 correponding to $\beta_{0,0},\beta_{2,0},\beta_{0,2}$, respectively, and $\bm{A}_{12}$ has only one nonzero entry $2a_{2,0}a_{1,1}+2a_{0,2}a_{1,1}$ correponding to $\gamma_{0,0}$. $\bm{A}_{21}$ and $\bm{A}_{22}$ are defined similarly.
We now look for the weights $\{\bm{\tilde{w}}^{k}\}_{k=1}^{K}$ satisfying (16) using the above intrinsic formulation (3.2.2). Notice from (17) that the coefficients $\bm{\beta}$ and $\bm{\gamma}$ can be solved from normal equations by regressions $\bm{\beta}=\bm{\Phi}^{{\dagger}}\mathbf{\tilde{u}}_{{\mathbf{x}_{0}}}^{1}$ and $\bm{\gamma}=\bm{\Phi}^{{\dagger}}\mathbf{\tilde{u}}_{{\mathbf{x}_{0}}}^{2}$, respectively, where $\bm{\Phi}^{{\dagger}}\in\mathbb{R}^{m\times K}$ is given in (4) and $\mathbf{\tilde{u}}_{{\mathbf{x}_{0}}}^{j}=(\tilde{u}^{j}({\mathbf{x}_{0,1}}),...,\tilde{u}^{j}(\mathbf{x}_{0,K}))^{\top}\in\mathbb{R}^{K\times 1}$ $(j=1,2)$. Then one can write $(\Delta_{B}u(\mathbf{x}_{0}))_{1}$ as:
$(\Delta_{B}u(\mathbf{x}_{0}))_{1}=\bm{A}_{11}\bm{\Phi}^{{\dagger}}\mathbf{\tilde{u}}_{{\mathbf{x}_{0}}}^{1}+\bm{A}_{12}\bm{\Phi}^{{\dagger}}\mathbf{\tilde{u}}_{{\mathbf{x}_{0}}}^{2}:=\bm{\tilde{w}}_{11}^{\bullet}\mathbf{\tilde{u}}_{{\mathbf{x}_{0}}}^{1}+\bm{\tilde{w}}_{12}^{\bullet}\mathbf{\tilde{u}}_{{\mathbf{x}_{0}}}^{2},$ (19)

### v2

@@ 3 GMLS using intrinsic formulation
In this section, we approximate differential operators for vector fields using intrinsic formulation. For 2D manifolds in $\mathbb{R}^{3}$, we first review GMLS for the local Monge patch parametrization of each stencil $S_{{\mathbf{x}_{0}}}$ [29, 22] in Section 3.1, and subsequently derive the numerical scheme to approximate the Bochner Laplacian acting on a vector field in Section 3.2. We then generalize our approach to approximating the Bochner Laplacian on $d$-dimensional manifolds embedded in $\mathbb{R}^{n}$ in Section 3.3, where we also provide a complexity analysis. In Section 3.4, we approximate the covariant derivative for 2D manifolds, noting that the formulation naturally generalizes to higher dimensions. For readers not familiar with intrinsic differential geometry, please refer to Section A.1 for a short review.
@@ 3.1 Monge parametrization
Then the “height” $q$ is approximated as a polynomial function of the local coordinates $(\theta_{1},\theta_{2})$ using the GMLS approach following the similar procedure introduced in Section 2.2. Here one key difference is the degrees of all “intrinsic” polynomials in the approximation space ranging from $2$ to $l$, that is, the approximation space is $\mathbb{P}_{\mathbf{x}_{0}}^{l}\backslash\mathbb{P}_{\mathbf{x}_{0}}^{1}=\mathrm{span}(\{\theta_{1}^{\alpha_{1}}\theta_{2}^{\alpha_{2}}\}_{2\leq\alpha_{1}+\alpha_{2}\leq l})$. Thus, the approximation of $q$ is the optimal solution to the least-squares problem (2):
@@
The local Riemannian metric $\bm{g}=(g_{ij})_{i,j=1}^{2}$ for ${\mathbf{x}}\in S_{{\mathbf{x}_{0}}}$ is approximated as
$\bm{g}=\left[\begin{matrix}1+\left(\partial_{\theta_{1}}\hat{q}\right)^{2}&\partial_{\theta_{1}}\hat{q}\partial_{\theta_{2}}\hat{q}\\ \partial_{\theta_{1}}\hat{q}\partial_{\theta_{2}}\hat{q}&1+\left(\partial_{\theta_{2}}\hat{q}\right)^{2}\end{matrix}\right].$ (12)
As usual, we use $g^{ij}$ to denote the inverse metric tensor.
@@ 3.2 Approximation of the Bochner Laplacian
In this section, we assume that $M$ is a two-dimensional manifold in $\mathbb{R}^{3}$. For any vector field given on the point cloud $\mathbf{X}_{M}\subset M$, we write it as $\bm{u}=u^{j}\bm{t}_{j}$ where the components $\{u^{1},u^{2}\}$ are defined on the nodes $\{\mathbf{x}_{i}\}_{i=1}^{N}$ and $\left\{\bm{t}_{1}(\mathbf{x}_{i}),\bm{t}_{2}(\mathbf{x}_{i})\right\}_{i=1}^{N}$ are global bases of tangent vector space at each node. Here and throughout this paper, we use the Einstein summation convention. Note that there is no need to do alignments of these bases (as shown in Fig. 1(a)). To obtain a finite-difference type approximation of the Bochner Laplacian at a base point $\mathbf{x}_{0}$, we intend to find the weights $\{\bm{w}^{k}\in\mathbb{R}^{2\times 2}\}_{k=1}^{K}$ such that,
$\Delta_{B}\bm{u}({\mathbf{x}_{0}})\approx\sum_{k=1}^{K}\bm{w}^{k}\bm{u}|_{{\mathbf{x}_{0,k}}}=(\bm{w}^{1},\ldots,\bm{w}^{K})\mathbf{u}_{{\mathbf{x}_{0}}},$ (13)
where $\{{\mathbf{x}_{0,k}}\}_{k=1}^{K}$ is the $K$-nearest neighbors of $\mathbf{x}_{0}$, $\bm{u}|_{{\mathbf{x}_{0,k}}}=(u^{1}({\mathbf{x}_{0,k}}),u^{2}({\mathbf{x}_{0,k}}))^{\top}$ and $\ \mathbf{u}_{{\mathbf{x}_{0}}}=(u^{1}({\mathbf{x}_{0,1}}),u^{2}({\mathbf{x}_{0,1}}),\ldots,u^{1}({\mathbf{x}_{0,K}}),u^{2}({\mathbf{x}_{0,K}}))^{\top}\in\mathbb{R}^{2K\times 1}.$
In Section 3.2.1, we derive the intrinsic formulation for the Bochner Laplacian in the local Monge coordinate system (11). In Section 3.2.2, we apply the GMLS to discretize the operator in the local Monge coordinate. In Section 3.2.3, we discuss how to obtain the weights in (13) and assemble these weights into a sparse $2N\times 2N$ matrix where $2$ corresponds to the intrinsic dimension of the surface.
@@ 3.2.1 Approximation of geometric quantities in the local Monge coordinate
The computations for geometric quantities and function derivatives involved in the Bochner Laplacian $\Delta_{B}\bm{u}$ could be complicated. Fortunately, here we only focus on the values of $\Delta_{B}\bm{u}$ at the base point ${\mathbf{x}_{0}}$ and their calculations can be much simplified in the Monge coordinate system as introduced in Section 3.1. In particular, we use the following properties of a Monge coordinate system:
@@
Let $\{U_{0};\xi^{1},\ldots,\xi^{d}\}$ be a Monge coordinate system at $\mathbf{x}_{0}$ where $U_{0}\subset M$ is an open neighborhood of $\mathbf{x}_{0}$ and $M$ is a $d$-dimensional manifold with Riemannian metric $\bm{g}$. Denote the Christoffel symbol by $\Gamma_{ij}^{k}$. Then
@@
The Bochner Laplacian acting on a vector field $\bm{u}=\tilde{u}^{i}\partial_{i}$ is defined as
$\Delta_{B}\bm{u}=g^{ij}\tilde{u}^{k}_{;ij}\partial_{k},$
@@
Applying Proposition 3.1 and noticing that the tangent plane bases in (11) are $\partial_{k}|_{\mathbf{x}_{0}}=\bm{t}_{k}({\mathbf{x}_{0}})$, one can simplify the Bochner Laplacian $\Delta_{B}\bm{u}(\mathbf{x}_{0})$ as,
$\Delta_{B}\bm{u}(\mathbf{x}_{0})=g^{ij}\tilde{u}_{;ij}^{k}\partial_{k}|_{\mathbf{x}_{0}}=\left[\left(\tilde{u}_{;11}^{1}+\tilde{u}_{;22}^{1}\right)\partial_{1}+\left(\tilde{u}_{;11}^{2}+\tilde{u}_{;22}^{2}\right)\partial_{2}\right]\big{|}_{\mathbf{x}_{0}},$
@@
where $a_{\alpha_{1},\alpha_{2}}$ are the coefficients of $\hat{q}$ as in (10). Similarly, one can obtain all the derivatives of Christoffel symbols at $\mathbf{x}_{0}$ for 2D manifolds in $\mathbb{R}^{3}$:
@@
where the restriction notations $|_{\mathbf{x}_{0}}$ are all omitted. Then one can calculate $\Delta_{B}\bm{u}(\mathbf{x}_{0})$ in (3.2.1) as
$\Delta_{B}\bm{u}(\mathbf{x}_{0})$ $=$ $\left(\partial_{11}\tilde{u}^{1}+\partial_{22}\tilde{u}^{1}+\tilde{u}^{1}\left[4(a_{2,0})^{2}+(a_{1,1})^{2}\right]+\tilde{u}^{2}\left[2a_{2,0}a_{1,1}+2a_{0,2}a_{1,1}\right]\right)\bm{t}_{1}({\mathbf{x}_{0}})$ (16)
@@
We notice that the Bochner Laplacian in (16) involves only the coefficients $\{a_{\alpha_{1},\alpha_{2}}\}$ that satisfy ${\alpha_{1}+\alpha_{2}=2}$, which correspond to the monomials in (10) of degree 2.
@@ 3.2.2 Approximation of function and its derivatives in the local Monge coordinate
Let $\bm{u}=\tilde{u}^{i}\partial_{i}$ be a vector field defined locally on the stencil $S_{\mathbf{x}_{0}}$, where $\left\{\partial_{1},\partial_{2}\right\}$ is the local basis defined in (11). Our goal here is to approximate the Bochner Laplacian $\Delta_{B}$ acting on the vector field $\bm{u}$ at ${\mathbf{x}_{0}}$ by a linear combination of the vector values on $S_{\mathbf{x}_{0}}$, i.e., $\{\bm{\tilde{u}}|_{{\mathbf{x}_{0,k}}}:=(\tilde{u}^{1}({\mathbf{x}_{0,k}}),\tilde{u}^{2}({\mathbf{x}_{0,k}}))^{\top}\in\mathbb{R}^{2\times 1}\}_{k=1}^{K}$. Thus, we are seeking the weights $\{\bm{\tilde{w}}^{k}\in\mathbb{R}^{2\times 2}\}_{k=1}^{K}$ such that,
$\Delta_{B}\bm{u}({\mathbf{x}_{0}})\approx\sum_{k=1}^{K}\bm{\tilde{w}}^{k}\bm{\tilde{u}}|_{{\mathbf{x}_{0,k}}}=(\bm{\tilde{w}}^{1},\ldots,\bm{\tilde{w}}^{K})\mathbf{\tilde{u}}_{{\mathbf{x}_{0}}}\in\mathbb{R}^{2\times 1},$ (17)
@@
The GMLS approximation of the vector field $\bm{u}=\tilde{u}^{i}\partial_{i}$ gives the GMLS solution to each component, i.e.,
@@
Substituting all above into $\Delta_{B}\bm{u}(\mathbf{x}_{0})$ in (16), one arrives at
$\Delta_{B}\bm{u}(\mathbf{x}_{0})=(\Delta_{B}\bm{u}(\mathbf{x}_{0}))_{1}\bm{t}_{1}({\mathbf{x}_{0}})+(\Delta_{B}\bm{u}(\mathbf{x}_{0}))_{2}\bm{t}_{2}({\mathbf{x}_{0}})$
@@
where the first component of $\Delta_{B}\bm{u}(\mathbf{x}_{0})$ is $(\Delta_{B}\bm{u}(\mathbf{x}_{0}))_{1}:=\bm{A}_{11}\bm{\beta}+\bm{A}_{12}\bm{\gamma}$ and the second component is $(\Delta_{B}\bm{u}(\mathbf{x}_{0}))_{2}:=\bm{A}_{21}\bm{\beta}+\bm{A}_{22}\bm{\gamma}$. Here $\bm{\beta}=\left(\beta_{\alpha_{1}(j),\alpha_{2}(j)}\right)_{j=1}^{m}\in\mathbb{R}^{m\times 1}$ and $\bm{\gamma}=\left(\gamma_{\alpha_{1}(j),\alpha_{2}(j)}\right)_{j=1}^{m}\in\mathbb{R}^{m\times 1}$ are both $m$ by $1$ column vectors and $\{\bm{A}_{ij}\in\mathbb{R}^{1\times m}\}_{i,j=1}^{2}$ are all $1$ by $m$ row vectors. Specifically, $\bm{A}_{11}$ has three nonzero entries $24(a_{2,0})^{2}+(a_{1,1})^{2},2,$ 2 correponding to $\beta_{0,0},\beta_{2,0},\beta_{0,2}$, respectively, and $\bm{A}_{12}$ has only one nonzero entry $2a_{2,0}a_{1,1}+2a_{0,2}a_{1,1}$ correponding to $\gamma_{0,0}$. $\bm{A}_{21}$ and $\bm{A}_{22}$ are defined similarly.
We now look for the weights $\{\bm{\tilde{w}}^{k}\}_{k=1}^{K}$ satisfying (17) using the above intrinsic formulation (3.2.2). Notice from (18) that the coefficients $\bm{\beta}$ and $\bm{\gamma}$ can be solved from normal equations by regressions $\bm{\beta}=\bm{\Phi}^{{\dagger}}\mathbf{\tilde{u}}_{{\mathbf{x}_{0}}}^{1}$ and $\bm{\gamma}=\bm{\Phi}^{{\dagger}}\mathbf{\tilde{u}}_{{\mathbf{x}_{0}}}^{2}$, respectively, where $\bm{\Phi}^{{\dagger}}\in\mathbb{R}^{m\times K}$ is given in (4) and $\mathbf{\tilde{u}}_{{\mathbf{x}_{0}}}^{j}=(\tilde{u}^{j}({\mathbf{x}_{0,1}}),...,\tilde{u}^{j}(\mathbf{x}_{0,K}))^{\top}\in\mathbb{R}^{K\times 1}$ $(j=1,2)$. Then one can write $(\Delta_{B}\bm{u}(\mathbf{x}_{0}))_{1}$ as:
$(\Delta_{B}\bm{u}(\mathbf{x}_{0}))_{1}=\bm{A}_{11}\bm{\Phi}^{{\dagger}}\mathbf{\tilde{u}}_{{\mathbf{x}_{0}}}^{1}+\bm{A}_{12}\bm{\Phi}^{{\dagger}}\mathbf{\tilde{u}}_{{\mathbf{x}_{0}}}^{2}:=\bm{\tilde{w}}_{11}^{\bullet}\mathbf{\tilde{u}}_{{\mathbf{x}_{0}}}^{1}+\bm{\tilde{w}}_{12}^{\bullet}\mathbf{\tilde{u}}_{{\mathbf{x}_{0}}}^{2},$ (20)